OpenStudy (anonymous):

If u(x)=4x-5, v(x)=x^2, and f(x)=1/x, find formulas for, u(v(f(x)))= u(f(v(x)))= v(u(f(x)))=

OpenStudy (anonymous):

So, u(v(f(x)))= 4x-5(x^2(1/x(x)))?

OpenStudy (stamp):

u(v(f(x))=\[\frac{4}{x^2}-5\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (stamp):

\[v(f(x))=(\frac{1}{x})^2\]

OpenStudy (stamp):

\[=\frac{1}{x^2}\]

OpenStudy (anonymous):

oh I was just plugging in the u and v I thought it needs to be solved.

OpenStudy (anonymous):

I did check and the answer in the back of the book and the answer you gave is right, but I am confused how you got it.

OpenStudy (anonymous):

Actually all I did was u(v(f(x)))= plug the u, v, and f, 4x-5(x^2(1/x(x)))? I didn't really know from here if I was doing it right.

OpenStudy (stamp):

\[u=4x-5,\ v=x^2,\ f=1/x\]\[u(v(f)),\ v(f)=(1/x)^2,\ u(1/x^2)=4(1/x^2)-5\]\[u(v(f))=4/x^2-5\]

OpenStudy (anonymous):

thanks @stamp that's clear